In an arithmetic sequence with the 2nd term 8 and the 6th term 20, find the 100th term?
Answer by richwmiller(17219)

an = a + (n - 1)d
20 = a + 5d,
8 = a + d,
12=4d
3=d
a=5
a100 = a + 99d
a100=297+5=302
a100=302